# Enhanced Rényi Entropy-Based Post-Quantum Key Agreement with Provable Security and Information-Theoretic Guarantees

### v1

@@ 4 Enhanced Theoretical Analysis
4-A Entropy Amplification Theorem
@@
Theorem 4 (Min-Entropy Preservation).
For independent random variables $X_{1},\dots,X_{n}$ over $\{0,1\}^{m}$ with $H_{\infty}(X_{i})\geq\gamma$, the XOR sum $S=\bigoplus_{i=1}^{n}X_{i}$ satisfies:
$H_{\infty}(S)\geq\max\left(0,n\gamma-(n-1)m\right)\geq\kappa$ (15)
@@
$\max_{z}\Pr[X\oplus Y=z]\leq\min\left(2^{-H_{\infty}(X)},2^{-H_{\infty}(Y)}\right)\cdot|\mathcal{X}|\cdot\alpha$ (17)
@@
$H_{\infty}(X\oplus Y)\geq\max\left(H_{\infty}(X),H_{\infty}(Y)\right)+\log|\mathcal{X}|-\beta$ (18)
@@
$\Pr[X\oplus Y=z]$
$\leq\max_{x}\Pr[X=x]\cdot\max_{y}\Pr[Y=y]$
$\leq 2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)}$ (20)
@@
$\geq\left[(n-1)\gamma-(n-2)\log|\mathcal{X}|\right]+\gamma-\log|\mathcal{X}|$
$=n\gamma-(n-1)\log|\mathcal{X}|.$ (23)
@@
Lemma 2 (Quantum Chain Rule for Min-Entropy).
@@
$S_{\infty}(S|E)\geq\sum_{i=1}^{n}S_{\infty}(X_{i}|X_{1}\cdots X_{i-1}E)-(n-1)\log|\mathcal{X}|.$ (25)
@@
4-B Security Parameterization Framework
Theorem 5 (Security Parameterization).
@@ 4-B Security Parameterization Framework
$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$ (26)
@@
$\delta(\rho_{KE},\tau_{K}\otimes\rho_{E})\leq\epsilon$ (27)
@@
$\quad m\geq\max(3\kappa,\gamma).$ (28)
@@
4-C Composable Security Framework
@@
Theorem 6 (Composable Security).
@@
  • Upon receiving $(\text{init})$ from all honest parties, output $K\xleftarrow{\$}\{0,1\}^{\kappa}$ to all parties
@@
    Commit phase: Generate random $c_{i}^{\prime}\xleftarrow{\$}\{0,1\}^{\lambda}$ without knowing $s_{i}$
@@
    Share phase: Simulate shares using random polynomials consistent with $c_{i}^{\prime}$
@@
    Verification phase: Program random oracle $\mathcal{H}$ to satisfy $\mathcal{H}(s_{i}^{\prime}\|\hat{H}_{i}^{\prime})=c_{i}^{\prime}$ for extracted $s_{i}^{\prime}$
@@
$\left|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{KA},\mathcal{S},\mathcal{Z}}=1]\right|\leq\mathsf{negl}(\kappa)$ (29)
@@ 5 Security Analysis
5-A Passive Security Against Quantum Eavesdroppers
Theorem 7 (Passive Security).
Under QROM, for any QPT passive adversary $\mathscr{A}$, the distinguishing advantage satisfies:
$\left|\Pr[\mathscr{A}(K)=1]-\Pr[\mathscr{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\mathsf{negl}(\lambda)$ (30)
@@ 5-A Passive Security Against Quantum Eavesdroppers
$\text{Adv}(\mathscr{A})$ $\leq\delta(\text{real protocol},\text{ideal protocol})$
$+\delta(\text{ideal protocol},\text{uniform}).$ (31)
@@
5-B Active Adversary Resistance
Theorem 8 (Active Security).
With authenticated channels, for any QPT active adversary $\mathscr{A}$:
$\Pr[\mathscr{A}\text{ wins}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}+\frac{n^{2}}{2^{\lambda}}+\mathsf{negl}(\kappa)$ (32)
@@ 5-B Active Adversary Resistance
Union bound gives the result. The $\frac{n^{2}}{2^{\lambda}}$ term accounts for commitment forgeries across all parties. ∎
5-C Quantum Attack Resilience Analysis
@@
$\gamma\geq\kappa+\log Q+\log(1/\epsilon).$ (33)
@@
$\gamma$ $=\left\lceil\frac{128+40+5\times 10+4\times 384}{5}\right\rceil$
$=\left\lceil\frac{128+40+50+1536}{5}\right\rceil=\lceil 350.8\rceil=351$
@@
Theorem 9 (Comprehensive Quantum Security).
For any QPT adversary $\mathscr{A}$ with quantum memory $Q=2^{q}$ qubits, the distinguishing advantage satisfies:
$\left|\Pr[\mathscr{A}(K)=1]-\Pr[\mathscr{A}(U)=1]\right|$
$\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\frac{n^{2}}{2^{\lambda}}+2^{q-(n\gamma-(n-1)m)}.$ (34)
@@
5-C1 Quantum Collision Resistance Analysis
The Brassard-Høyer-Tapp (BHT) attack [7] achieves $O(2^{m/3})$ complexity by:
[…]
$\text{Measure collision with prob. }p\geq c\cdot 2^{-m/3}.$
@@ 5-C1 Quantum Collision Resistance Analysis
$\text{Expected queries}=\sqrt{\frac{\pi}{4p}}\geq\sqrt{\frac{\pi}{4c}}\cdot 2^{\kappa}\gg 2^{\kappa}.$ (35)
@@ 6 Entropy Requirements and Parameter Trade-offs
6-A Theoretical Parameter Optimization
@@
6-A1 Quantum-Secure Entropy Estimation
@@
Theorem 10 (Quantum Min-Entropy Sampling).
For $\epsilon>0$ and samples $\mathcal{S}=\{x_{1},\dots,x_{N}\}$ from source $X$:
$\Pr_{x_{i}\leftarrow X}\left[\hat{H}_{\infty}(X)\geq S_{\infty}(X|E)_{\rho}-\Delta\right]\geq 1-\epsilon$ (37)
@@
6-B Quantum Advantage Mitigation
@@
6-B1 Quantum Effects on Entropy Estimation
@@
$\sup_{\rho}\left|\hat{H}_{\infty}^{\text{class}}(X)-S_{\infty}(X|E)_{\rho}\right|\leq\delta+\log(1/\epsilon)$ (38)
@@
    Quantum-secure randomness extractors: $\text{Ext}:\{0,1\}^{m}\times\{0,1\}^{d}\rightarrow\{0,1\}^{\kappa}$
@@
    $\Pr_{x^{n}\leftarrow X^{n}}\left[\hat{H}_{\infty}(X)\geq H_{\infty}(X)-\Delta\right]\geq 1-\epsilon$ (39)
@@ 7 Comparative Analysis and Protocol Extensions
7-A Comparative Analysis with Post-Quantum Alternatives
The quantum security landscape features diverse approaches with fundamentally different security foundations. Our protocol’s distinctive information-theoretic security provides unique advantages compared to computational post-quantum solutions. As shown in Table 3, while lattice-based schemes like CRYSTALS-Kyber offer efficient $\mathcal{O}(n\kappa)$ communication, their security relies on the unproven hardness of module-LWE problems, which remain vulnerable to unforeseen quantum algorithmic breakthroughs. Similarly, isogeny-based schemes like SIKE provide compact key sizes but have suffered devastating cryptanalytic attacks in recent years, demonstrating the fragility of dependency on specific mathematical assumptions.
TABLE 3: Post-Quantum Key Agreement Protocol Comparison
Protocol Security Basis Sec. Type Comm. Comp. Hardware Trust Hardness
Our Protocol Rényi entropy Info-theoretic $\mathcal{O}(n^{2}m)$ $\mathcal{O}(n^{2})$ Classical Honest majority None
CRYSTALS-Kyber [8] Module-LWE Computational $\mathcal{O}(n\kappa)$ $\mathcal{O}(n\kappa^{2})$ Classical PKI MLWE
SPHINCS+ [28] Hash functions QROM $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa)$ Classical None QROM
SIKE [11] Isogenies Computational $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa^{4})$ Classical PKI SSI
QKD [25] Quantum physics Info-theoretic $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa)$ Quantum Auth. chan. None
Sec. Type: Security Type; Comm.: Communication Complexity; Comp.: Computational Complexity;
Hardware: Hardware Requirements; Trust: Trust Assumptions; Hardness: Hardness Assumptions.
Auth. chan.: Authenticated channels; PKI: Public Key Infrastructure.
@@ 7-A Comparative Analysis with Post-Quantum Alternatives
7-B Protocol Extensions to Secure Applications
@@
1. Secure Multiparty Computation for Linear Functions: The protocol directly supports privacy-preserving computation of linear functions $f(s_{1},\dots,s_{n})=\sum c_{i}s_{i}$ without revealing individual inputs. This enables:
@@
The derived key $K_{f}=\mathcal{H}(f(s_{1},\dots,s_{n}))$ maintains the min-entropy guarantee $H_{\infty}(K_{f})\geq\kappa$ when coefficients $c_{i}$ are properly constrained.
2. Quantum-Hybrid Security Architecture: Integrating with quantum key distribution creates a defense-in-depth architecture:
$\text{Hybrid Key}=\mathcal{H}(K_{\text{QKD}}\oplus K_{\text{Entropy}})$ (40)
@@
3. Post-Quantum Authentication Framework: We develop an entropy-based message authentication code (MAC):
$\text{MAC}(k,m,s)=\mathcal{H}(k\oplus\mathcal{H}(m\|s))$ (41)
leveraging the same entropy sources used in key agreement. This provides quantum-resistant authentication with security bound $\Pr[\text{forge}]\leq\frac{q^{2}}{2^{\gamma}}+\frac{(q+1)^{2}}{2^{\lambda}}$, integrating seamlessly with our key establishment protocol.
7-C Future Research Directions
@@
1) Lightweight Implementations for IoT: Developing optimized implementations for resource-constrained devices presents significant challenges. The quadratic communication complexity $\mathcal{O}(n^{2}m)$ becomes problematic for large $n$, requiring compression techniques for shares and commitments. Research should explore:
@@
2) Blockchain Integration for Randomness Beacons: The protocol can power decentralized randomness generation for blockchain consensus:
$K_{\text{epoch}}=\mathcal{H}\left(\bigoplus_{i=1}^{n}\mathcal{H}(s_{i}\|\text{block}_{h-1})\right)$ (42)
@@
3) Post-Quantum Cryptography Standardization: Our work provides foundations for standardization efforts in information-theoretic PQC. Key initiatives include:
@@
  • Implementation testing frameworks Side-channel resistance certification
4) Entropy Source Diversity and Robustness: Future work must address practical challenges in entropy source management:
@@ 8 Conclusion
We have presented a theoretical framework for post-quantum key agreement based on Rényi entropy. The enhanced protocol addresses vulnerabilities in prior constructions through innovations: (1) a confidentiality-preserving verification mechanism using distributed polynomial commitments, (2) provably non-negative min-entropy bounds for XOR composition, and (3) composable security proofs in the quantum universal composability model.
@@
Entropy Preservation: $H_{\infty}(S)$ $\geq n\gamma-(n-1)m$ (43)
Verification Security: $\Pr[\text{bypass}]$ $\leq 2^{-\kappa}$ (44)
Composable Security: $\delta_{\text{QUC}}$ $\leq\operatorname{negl}(\kappa)$ (45)
Equations (43)-(45) form an integrated security framework that resists quantum attacks through fundamental information-theoretic principles rather than computational assumptions. For 128-bit security, our parameter optimization yields:
@@
Biography Section
Ruopengyu Xu received his B.S degree in network engineering from Software Engineering Institute Gungzhou in June 2025. He is currently an independent scientific researcher. Recently, he has published a paper in https://arxiv.org/abs/2506.07640. At the same time, he is also preparing his postgraduate plan. His research interests are in Quantum cryptography, Information Security, Number Theory and Public key cryptosystem so on.
Chenglian Liu received his B.S. degree in information management from National Union University in 1992, and his M.Sc. degree in National Defense from National Defense University in 2004. He was a research fellow in the Department of Automation Engineering at Technological Educational Institution of Sterea Hellas from 2017 to 2019, and a visiting scholar at Universidad del Pacifico from 2020 to 2022. He is currently an associate professor at the School of Electrical and Computer Engineering, Nanfang University, Guangzhou, China. His main research focuses on Information Security, Cryptanalysis, Fuzzy Sets, and Goldbach’s Conjecture, among other topics.

### v2

@@ 4 Enhanced Theoretical Analysis
4.1 Entropy Amplification Theorem
@@
Theorem 4.1 (Min-Entropy Preservation).
For independent random variables $X_{1},\ldots,X_{n}$ over $\{0,1\}^{m}$ with $H_{\infty}(X_{i})\geq\gamma$, the XOR sum $S=\bigoplus_{i=1}^{n}X_{i}$ satisfies:
$H_{\infty}(S)\geq\max\left(0,n\gamma-(n-1)m\right)\geq\kappa$
@@
$\max_{z}\Pr[X\oplus Y=z]\leq\min\left(2^{-H_{\infty}(X)},2^{-H_{\infty}(Y)}\right)\cdot|\mathcal{X}|\cdot\alpha$
@@
$H_{\infty}(X\oplus Y)\geq\max\left(H_{\infty}(X),H_{\infty}(Y)\right)+\log|\mathcal{X}|-\beta$
@@
$\Pr[X\oplus Y=z]$ $\leq\max_{x}\Pr[X=x]\cdot\max_{y}\Pr[Y=y]$
$\leq 2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)}$
@@
$\geq[(n-1)\gamma-(n-2)\log|\mathcal{X}|]+\gamma-\log|\mathcal{X}|$
$=n\gamma-(n-1)\log|\mathcal{X}|.$
@@
Lemma 4.2 (Quantum Chain Rule for Min-Entropy).
@@
$S_{\infty}(S|E)\geq\sum_{i=1}^{n}S_{\infty}(X_{i}|X_{1}\cdots X_{i-1}E)-(n-1)\log|\mathcal{X}|.$
@@
4.2 Security Parameterization Framework
Theorem 4.4 (Security Parameterization).
@@ 4.2 Security Parameterization Framework
$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$
@@
$\delta(\rho_{KE},\tau_{K}\otimes\rho_{E})\leq\epsilon$
@@
$\quad m\geq\max(3\kappa,\gamma)$
@@
4.3 Composable Security Framework
@@
Theorem 4.5 (Composable Security).
@@
  • 1.
    Upon receiving (init) from all honest parties, output $K\overset{\mathdollar}{\leftarrow}\{0,1\}^{\kappa}$ to all parties
  • 2.
@@
    Commit phase: Generate random $c^{\prime}_{i}\overset{\mathdollar}{\leftarrow}\{0,1\}^{\lambda}$ without knowing $s_{i}$
@@
    Share phase: Simulate shares using random polynomials consistent with $c^{\prime}_{i}$
@@
    Verification phase: Program random oracle $\mathcal{H}$ to satisfy $\mathcal{H}(s^{\prime}_{i}||\bar{H}^{\prime}_{i})=c^{\prime}_{i}$ for extracted $s^{\prime}_{i}$
@@
$|\Pr[\textsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\textsf{IDEAL}_{\mathcal{F}_{KA},\mathcal{S},\mathcal{Z}}=1]|\leq\textsf{negl}(\kappa)$
@@ 5 Security Analysis
5.1 Passive Security Against Quantum Eavesdroppers
Theorem 5.1 (Passive Security).
Under QROM, for any QPT passive adversary $\mathcal{A}$, the distinguishing advantage satisfies:
$\left|\Pr[\mathcal{A}(K)=1]-\Pr[\mathcal{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\mathsf{negl}(\lambda)$
@@ 5.1 Passive Security Against Quantum Eavesdroppers
$\mathsf{Adv}(\mathcal{A})\leq\delta(\text{real protocol},\text{ideal protocol})+\delta(\text{ideal protocol},\text{uniform}).$
@@
5.2 Active Adversary Resistance
Theorem 5.2 (Active Security).
With authenticated channels, for any QPT active adversary $\mathcal{A}$:
$\Pr[\mathcal{A}\text{ wins}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}+\frac{n^{2}}{2^{\lambda}}+\mathsf{negl}(\kappa)$
@@ 5.2 Active Adversary Resistance
Union bound gives the result. The $\frac{n^{2}}{2^{\kappa}}$ term accounts for commitment forgeries across all parties. ∎
5.3 Quantum Attack Resilience Analysis
@@
$\gamma\geq\kappa+\log Q+\log(1/\epsilon).$
@@
$\gamma$ $=\frac{128+40+5\times 10+4\times 384}{5}$
$=\frac{128+40+50+1536}{5}=\frac{1754}{5}=350.8\approx 351$
@@
Theorem 5.3 (Comprehensive Quantum Security).
For any QPT adversary $\mathcal{A}$ with quantum memory $Q=2^{q}$ qubits, the distinguishing advantage satisfies:
$\left|\Pr[\mathcal{A}(K)=1]-\Pr[\mathcal{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\frac{n^{2}}{2^{\lambda}}+2^{q-(n\gamma-(n-1)m)}.$
@@
5.3.1 Quantum Collision Resistance Analysis
The Brassard-Hoyer-Tapp (BHT) attack [7] achieves $O(2^{m/3})$ complexity by:
Phase 1: Create $\frac{\pi}{4}2^{m/3}$ states $|\psi_{i}\rangle=\sum\limits_{x}\alpha_{x}|x\rangle$
Phase 2: Apply $U_{f}:|x\rangle\rightarrow(-1)^{f(x)}|x\rangle$
Measure collision with prob. $p\geq c\cdot 2^{-m/3}$.
@@ 5.3.1 Quantum Collision Resistance Analysis
$\text{Expected queries}=\sqrt{\frac{\pi}{4p}}\geq\sqrt{\frac{\pi}{4c}}\cdot 2^{\kappa}\gg 2^{\kappa}.$
@@ 6 Entropy Requirements and Parameter Trade-offs
6.1 Theoretical Parameter Optimization
@@
6.1.1 Quantum-Secure Entropy Estimation
@@
Theorem 6.1 (Quantum Min-Entropy Sampling).
For $\epsilon>0$ and samples $\mathcal{S}=\{x_{1},\ldots,x_{N}\}$ from source $X$:
$\Pr_{x_{i}\leftarrow X}\left[\hat{H}_{\infty}(X)\geq S_{\infty}(X|E)_{\rho}-\Delta\right]\geq 1-\epsilon$
@@
6.2 Quantum Advantage Mitigation
@@
6.2.1 Quantum Effects on Entropy Estimation
@@
$\sup_{\rho}\left|\hat{H}^{\text{class}}_{\infty}(X)-S_{\infty}(X|E)_{\rho}\right|\leq\delta+\log(1/\epsilon)$
@@
    Quantum-secure randomness extractors: Ext : $\{0,1\}^{m}\times\{0,1\}^{d}\rightarrow\{0,1\}^{\kappa}$
@@
    $\Pr_{x^{n}\leftarrow X^{n}}\left[\hat{H}_{\infty}(X)\geq H_{\infty}(X)-\Delta\right]\geq 1-\epsilon$
@@ 7 Comparative Analysis and Protocol Extensions
7.1 Comparative Analysis with Post-Quantum Alternatives
The quantum security landscape features diverse approaches with fundamentally different security foundations. Our protocol’s distinctive information-theoretic security provides unique advantages compared to computational post-quantum solutions. As shown in Table 2, while lattice-based schemes like CRYSTALS-Kyber [8] offer efficient $\mathcal{O}(n\kappa)$ communication, their security relies on the unproven hardness of module-LWE problems, which remain vulnerable to unforeseen quantum algorithmic breakthroughs. Similarly, isogeny-based schemes like SIKE [11] provide compact key sizes but have suffered devastating cryptanalytic attacks in recent years, demonstrating the fragility of dependency on specific mathematical assumptions.
@@ 7.1 Comparative Analysis with Post-Quantum Alternatives
Table 2: Comparison with Post-Quantum Alternatives
Approach Security Basis Advantages Limitations
Our Protocol Information-theoretic Unconditional security, Quantum resistance, Fault tolerance Quadratic communication, Entropy source requirement
Lattice-based (e.g., Kyber) Computational (LWE) Efficient, Standardized Vulnerable to quantum algorithmic advances
Code-based Computational (Decoding) Mature theory, Conservative security Large key sizes, Not efficient
QKD Information-theoretic Proven security, Commercial availability Requires quantum channels, Distance limitations
7.2 Protocol Extensions to Secure Applications
@@
  1. 1.
    Secure Multiparty Computation for Linear Functions: The protocol directly supports privacy-preserving computation of linear functions $f(s_{1},\ldots,s_{n})=\sum c_{i}s_{i}$ without revealing individual inputs. This enables:
    • (a)
@@
    The derived key $K_{f}=\mathcal{H}(f(s_{1},\ldots,s_{n}))$ maintains the min-entropy guarantee $H_{\infty}(K_{f})\geq\kappa$ when coefficients $c_{i}$ are properly constrained.
  2. 2.
    Quantum-Hybrid Security Architecture: Integrating with quantum key distribution creates a defense-in-depth architecture:
    $\text{Hybrid Key}=\mathcal{H}(K_{\text{QKD}}\oplus K_{\text{Entropy}})$
@@
  3. 3.
    Post-Quantum Authentication Framework: We develop an entropy-based message authentication code (MAC):
    $\text{MAC}(k,m,s)=\mathcal{H}(k\oplus\mathcal{H}(m\|s))$
    leveraging the same entropy sources used in key agreement. This provides quantum-resistant authentication with security bound $\Pr[\text{forge}]\leq\frac{q^{2}}{2\gamma}+\frac{(q+1)^{2}}{2\lambda}$, integrating seamlessly with our key establishment protocol.
7.3 Future Research Directions
@@
  1. 1.
    Lightweight Implementations for IoT: Developing optimized implementations for resource-constrained devices presents significant challenges. The quadratic communication complexity $\mathcal{O}(n^{2}m)$ becomes problematic for large $n$, requiring compression techniques for shares and commitments. Research should explore:
    • (a)
@@
  2. 2.
    Blockchain Integration for Randomness Beacons: The protocol can power decentralized randomness generation for blockchain consensus:
    $K_{\text{epoch}}=\mathcal{H}\left(\bigoplus_{i=1}^{n}\mathcal{H}(s_{i}\|\text{block}_{h-1})\right)$
@@
  3. 3.
    Post-Quantum Cryptography Standardization: Our work provides foundations for standardization efforts in information-theoretic PQC. Key initiatives include:
    • (a)
@@
    • (c)
      Implementation testing frameworks
    • (d)
      Side-channel resistance certification
  4. 4.
    Entropy Source Diversity and Robustness: Future work must address practical challenges in entropy source management:
    • (a)
@@ 8 Conclusion
We have presented a theoretical framework for post-quantum key agreement based on Renyi entropy. The enhanced protocol addresses vulnerabilities in prior constructions through innovations: (1) a confidentiality-preserving verification mechanism using distributed polynomial commitments, (2) provably non-negative min-entropy bounds for XOR composition, and (3) composable security proofs in the quantum universal composability model.
@@
Entropy Preservation: $\quad H_{\infty}(S)\geq n\gamma-(n-1)m$
Verification Security: $\quad\Pr[\text{bypass}]\leq 2^{-\kappa}$
Composable Security: $\quad\delta_{\text{QUC}}\leq\text{negl}(\kappa)$
These equations form an integrated security framework that resists quantum attacks through fundamental information-theoretic principles rather than computational assumptions. For 128-bit security, our parameter optimization yields:
@@
Appendix A Proofs of Technical Lemmas
This appendix contains detailed proofs of selected technical results from the main text.
A.1 Proof of Theorem 4 (Complete Version)
The complete proof of the min-entropy preservation theorem involves careful analysis of the convolution properties of probability distributions over finite fields…